OpenStudy (mayankdevnani):

From the divisors of 1980,how many of them are divisible by 4 but not by 15. (Question from PERMUTATION AND COMBINATION)

OpenStudy (zarkon):

find the prime factorization...then the problem is pretty easy

OpenStudy (mayankdevnani):

prime factorization of 1980=\(\large \bf 2^2 \times 3^2 \times 5 \times 11\)